OpenStudy (anonymous):

Is the answer -9x? Find the difference. 8x - (12x2 + 5x) =

OpenStudy (anonymous):

Help! haha

OpenStudy (konradzuse):

are their choices on this one, not sure exactly what they are referring to for difference... Unless of 2 squares?

OpenStudy (anonymous):

Yeah there is

OpenStudy (konradzuse):

Or what they want us to find.... Ok, can you post the choices please?

OpenStudy (valpey):

Are you simplifying: \[8x-(12x^2+5x)=\]

OpenStudy (anonymous):

here it is A. -9xB. -12x2 + 3xC. -12x2 + 5xD. 12x2 + 3x

OpenStudy (valpey):

B

OpenStudy (konradzuse):

ok so it would just be simplifying then... 8x-12x^2 +5x

OpenStudy (konradzuse):

8x-5x = 3x

OpenStudy (konradzuse):

-12x^2 +3x so B

OpenStudy (valpey):

\[8x-(12x^2+5x)=8x-12x^2-5x=-12x^2+8x-5x=-12x^2+3x\]

OpenStudy (anonymous):

ohh haha i didnt htink of that haha i have never seen this before so thanks guys!

OpenStudy (konradzuse):

yup.. Idk why it's asking for a difference lol...

OpenStudy (konradzuse):

oh no medal this time :( :p